OpenStudy (anonymous):

Help please a company uses two vans to transport workers from a free parking lot to the workplace between 7:00 and 9:00a.m. one van has 6 more seats than the other. the smaller van makes two trips every morning while the larger one makes only one trip. the two vans can transport 57 people, maximum. Let x be the seats in the small van while y be the seats for the larger van. how many seats does the larger van have?

OpenStudy (anonymous):

Let the number of seats in the small van = x Then the number of seats in the big van = x+6

OpenStudy (anonymous):

if they can together transport 57 people a day then set your equation equal to 57

OpenStudy (anonymous):

since the small van takes two trips then the number of people transported each day is 2x

OpenStudy (anonymous):

You add both the vans together and set it equal to 57

OpenStudy (anonymous):

2x+x+6=57 3x+6=57 then you solve for x